# Neural Point Cloud Diffusion for Disentangled 3D Shape and Appearance Generation

### v1

@@ Abstract
Controllable generation of 3D assets is important for many practical applications like content creation in movies, games and engineering, as well as in AR/VR. Recently, diffusion models have shown remarkable results in generation quality of 3D objects. However, none of the existing models enable disentangled generation to control the shape and appearance separately. For the first time, we present a suitable representation for 3D diffusion models to enable such disentanglement by introducing a hybrid point cloud and neural radiance field approach. We model a diffusion process over point positions jointly with a high-dimensional feature space for a local density and radiance decoder. While the point positions represent the coarse shape of the object, the point features allow modeling the geometry and appearance details. This disentanglement enables us to sample both independently and therefore to control both separately. Our approach sets a new state of the art in generation compared to previous disentanglement-capable methods by reduced FID scores of 30-90% and is on-par with other non-disentanglement-capable state-of-the art methods. Code is provided at https://github.com/lmb-freiburg/neural-point-cloud-diffusion.
@@ 1 Introduction
To establish our model, we first train a generalizable Point-NeRF renderer [42] by sharing its weights across many instances of ShapeNet [36] or PhotoShape [29] objects. The obtained neural point clouds then serve as a dataset to train a diffusion model that learns to denoise the point positions and features simultaneously. Different to previous diffusion models, our model operates on high-dimensional latent spaces. In summary, our contributions are:
@@ 2 Related work
NeRF [24] represents geometry and appearance as a radiance and density field that can be volumetrically rendered to photorealistic images. Point-NeRF [42] extends NeRF to a parameterization by a point cloud that is obtained from MVSNet [44]. This reduces ambiguity and allows for a much faster reconstruction. The Point-NeRF MLPs are originally trained on a single scene. In this work, we train them jointly on many objects to obtain a generalizable version [41]. In addition, we regularize the neural point cloud features to be optimally suitable for the diffusion model, which we use to generate objects.
@@ 3 Method
In this section, we describe Neural Point Cloud Diffusion (NPCD), our generative model for 3D shape and appearance via diffusion on neural point clouds. An overview of the method is shown in Fig. 1. At the center of our method is a generalizable neural point representation, which is further described in Sec. 3.1. We discuss characteristics of autodecoder schemes in Sec. 3.2 and provide regularization schemes that enable denoising diffusion on the feature space. Subsequently, in Sec. 3.3 we then present a diffusion model to denoise the neural point positions and features. After the diffusion model is trained, we can sample 3D shape and appearance independently from each other as described in Sec. 3.4.
@@
3.1 Generalizable Point-NeRF
@@
with $\mathcal{L}$ being the mean squared error between rendered and ground truth pixel colors. In contrast to Point-NeRF, we share parameters $\phi,\psi,\gamma$ of the rendering MLPs over all objects, obtaining a generalizable representation.
@@ 3.3 Neural point cloud diffusion
Diffusion models [37, 38] learn the distribution $q(\mathbf{x})$ of data $\mathbf{x}$ by defining a forward diffusion process in form of a Markov chain with steps $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ that gradually transform the data distribution into a simple known distribution. A model $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ with parameters $\theta$ is then trained to approximate the steps $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ in the Markov chain of the reverse process, which allows to evaluate the likelihood of a given data point, or sample from the learnt distribution.
@@
The distribution of both modalities, point positions $\mathbf{P}$ and appearance features $\mathbf{F}$, is learnt jointly. During training, we sample Gaussian noise $\mathbf{\epsilon}^{\mathbf{P}}$ for all point positions and $\mathbf{\epsilon}^{\mathbf{F}}$ for all point features and use it to obtain the noised neural point cloud $\mathcal{P}_{t}=(\mathbf{P}_{t},\mathbf{F}_{t})$ at a specific timestep $t$ in the diffusion process via Eq. 10. The denoiser network $T_{\theta}((\mathbf{P}_{t},\mathbf{F}_{t}),t)=(\mathbf{\epsilon}^{\mathbf{P}}_{\theta},\mathbf{\epsilon}^{\mathbf{F}}_{\theta})$ takes the noised neural point cloud and timestep as input and estimates the noise $\mathbf{\epsilon}^{\mathbf{P}}_{\theta}$ and $\mathbf{\epsilon}^{\mathbf{F}}_{\theta}$ that was applied to the points and features. The network is optimized by minimizing the average mean squared error on both noise vectors:
$\mathcal{L}_{T}=\frac{1}{2}(\text{MSE}(\mathbf{\epsilon}^{\mathbf{P}},\mathbf{\epsilon}^{\mathbf{P}}_{\theta})+\text{MSE}(\mathbf{\epsilon}^{\mathbf{F}},\mathbf{\epsilon}^{\mathbf{F}}_{\theta}))\,.$ (11)
@@
As the architecture for the denoiser network we use a Transformer [26, 40]. As input, the transformer receives $M+1$ tokens, one token per point plus one additional token encoding $t$. Point tokens are obtained by concatenating the point position and feature of each point in the noisy point cloud and projecting them with a linear layer. Similarly, the $t$ token is obtained by projection with its own linear layer onto the same dimensionality. After encoding, all tokens are processed by transformer layers, including self-attention and MLPs. Finally, the resulting output tokens corresponding to the $M$ points are projected back to the dimensions of the input point positions and features and interpreted as noise predictions $\mathbf{\epsilon}^{\mathbf{P}}_{\theta}$ and $\mathbf{\epsilon}^{\mathbf{F}}_{\theta}$.
@@ 3.4 Disentangled generation
To sample from $p(\mathbf{F}|\mathbf{P})$ given point positions $\mathbf{P}_{0}$, we obtain the initial noisy neural point cloud $\mathcal{P}_{T}=(\mathbf{P}_{T},\mathbf{F}_{T})$ by
$\mathbf{P}_{T}=\sqrt{\bar{\alpha}_{t}}\mathbf{P}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon}^{\mathbf{P}},\quad\mathbf{F}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\textnormal{,}$ (12)
i.e., only sampling $\mathbf{F}_{T}$ from noise while obtaining $\mathbf{P}_{T}$ as a noisy variant of $\mathbf{P}_{0}$. Then, the denoising transformer is applied. After each reverse process iteration, we reset $\mathbf{P}_{t}$ to the values from Eq. 12 for the appropriate $t$. Sampling from $p(\mathbf{P}|\mathbf{F})$ can be done analogously.
Note that the conditional sampling procedure described above is enabled by the neural point representations, which allow to individually modify the disentangled variables for coarse shape and local appearance.
@@ 4.1 Datasets and experimental setup
Besides ShapeNet SRN, we also use the PhotoShape Chairs dataset [29]. The dataset contains $15{,}576$ objects and features more realistic textures on top of ShapeNet meshes. We use the same test split as DiffRF [25], which consists of $1,552$ objects. From the remaining objects, we randomly select $2,480$ objects for training. Further, we use the same renderings as DiffRF [25], which consist of 200 views per object on an Archimedean spiral. We use images at a resolution of 128x128 pixels. We use the same point clouds with 512 points as for SRN chairs.
@@
Figure 2: Qualitative examples of disentangled generation on SRN cars, SRN chairs, PhotoShape chairs. 1(a) Appearance-only generation: we show a generated object and objects with re-sampled appearance. 1(b) Shape-only generation: we show a generated object and objects with re-sampled coarse shape. We can get diverse samples of local appearance or coarse shape when the respective other is given.
@@
We construct neural point clouds by zero-initializing features for each point. In generalizable Point-NeRF, we optimize the reconstruction loss in Eq. 3 with the TV and KL regularizers from Eq. 4 and Eq. 5. Further details on network architectures and training parameters are provided in the supplementals. For the diffusion model training, we normalize the neural point clouds and use DDPM [13] diffusion model parameters. Further details on the denoiser architecture, diffusion model parameters, and training parameters are provided in the supplementals.
@@ 4.6 Analysis
As diffusion on hybrid point clouds and local radiance fields has not been done before, we conduct ablation studies and analyze various novel design choices. Here, we analyze the effects of different initialization strategies, feature dimensionality and regularization methods in the generalizable Point-NeRF and diffusion model. We provide a more detailed analysis in the supplementals.
@@
As we assume that the structure of the feature space strongly affects the diffusion model, we analyze the effects of applying KL and TV regularizations to the neural point features during the generalizable Point-NeRF training. Tab. 4 shows that both regularizations further decrease the ambiguity of the latent space over the zero initialization. We can summarize that a combination of TV and KL regularization provides overall the best results (c.f. supplementals). Overall, we found an appropriate initialization and regularization to be key ingredients for successful neural point cloud diffusion.

### v2

@@ Abstract
Controllable generation of 3D assets is important for many practical applications like content creation in movies, games and engineering, as well as in AR/VR. Recently, diffusion models have shown remarkable results in generation quality of 3D objects. However, none of the existing models enable disentangled generation to control the shape and appearance separately. For the first time, we present a suitable representation for 3D diffusion models to enable such disentanglement by introducing a hybrid point cloud and neural radiance field approach. We model a diffusion process over point positions jointly with a high-dimensional feature space for a local density and radiance decoder. While the point positions represent the coarse shape of the object, the point features allow modeling the geometry and appearance details. This disentanglement enables us to sample both independently and therefore to control both separately. Our approach sets a new state of the art in generation compared to previous disentanglement-capable methods by reduced FID scores of 30-90% and is on-par with other non-disentanglement-capable state-of-the art methods.
@@ 1 Introduction
To establish our model, we first train a category-level Point-NeRF autodecoder [50] by sharing its weights across many instances of ShapeNet [44] or PhotoShape [34] objects. The obtained neural point clouds then serve as a dataset to train a diffusion model that learns to denoise the point positions and features simultaneously. Different to previous diffusion models, our model operates on high-dimensional latent spaces. In summary, our contributions are:
@@ 2 Related work
NeRF [29] represents geometry and appearance as a radiance and density field that can be volumetrically rendered to photorealistic images. Point-NeRF [50] extends NeRF to a parameterization by a point cloud that is obtained from MVSNet [52]. This reduces ambiguity and allows for a much faster reconstruction. The Point-NeRF MLPs are originally trained on a single scene. In this work, we train them jointly on many objects as a category-level autodecoder [49]. In addition, we regularize the neural point cloud features to be optimally suitable for the diffusion model, which we use to generate objects.
@@ 3 Method
In this section, we describe Neural Point Cloud Diffusion (NPCD), our generative model for 3D shape and appearance via diffusion on neural point clouds. An overview of the method is shown in Fig. 1. At the center of our method is an autodecoder with a neural point representation for the latent codes, which is further described in Sec. 3.1. We discuss characteristics of autodecoder schemes in Sec. 3.2 and provide regularization schemes that enable denoising diffusion on the feature space. Subsequently, in Sec. 3.3 we then present a diffusion model to denoise the neural point positions and features. After the diffusion model is trained, we can sample 3D shape and appearance independently from each other as described in Sec. 3.4. We provide code for our method at https://github.com/lmb-freiburg/neural-point-cloud-diffusion.
@@
3.1 Category-Level Point-NeRF Autodecoder
@@
with $\mathcal{L}$ being the mean squared error between rendered and ground truth pixel colors. In contrast to Point-NeRF, we share parameters $\phi,\psi,\gamma$ of the rendering MLPs over all objects as an autodecoder with $(\mathbf{P}_{j},\mathbf{F}_{j})$ as latent codes.
@@ 3.3 Neural point cloud diffusion
Diffusion models [45, 46] learn the distribution $q(\mathbf{x})$ of data $\mathbf{x}$ by defining a forward diffusion process in form of a Markov chain with steps $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ that gradually transform the data distribution into a simple known distribution. A model $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ with parameters $\theta$ is then trained to approximate the steps $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ in the Markov chain of the reverse process, which allows to sample from the learnt distribution.
@@
The distribution of both modalities, point positions $\mathbf{P}$ and appearance features $\mathbf{F}$, is learnt jointly. During training, we sample Gaussian noise $\mathbf{\epsilon}^{\mathbf{P}}$ for all point positions and $\mathbf{\epsilon}^{\mathbf{F}}$ for all point features and use it to obtain the noised neural point cloud $\mathcal{P}_{t}=(\mathbf{P}_{t},\mathbf{F}_{t})$ at a specific timestep $t$ in the diffusion process via Eq. 10. The denoiser network $T_{\theta}((\mathbf{P}_{t},\mathbf{F}_{t}),t)=(\mathbf{\epsilon}^{\mathbf{P}}_{\theta},\mathbf{\epsilon}^{\mathbf{F}}_{\theta})$ takes the noised neural point cloud and timestep as input and estimates the noise $\mathbf{\epsilon}^{\mathbf{P}}_{\theta}$ and $\mathbf{\epsilon}^{\mathbf{F}}_{\theta}$ that was applied to the points and features. The network is optimized by minimizing the average mean squared error on both noise vectors.
@@
As architecture for the denoiser network, we use a Transformer [31, 48, 35]. As input, the transformer receives $M+1$ tokens, one token per point plus one additional token encoding $t$. Point tokens are obtained by concatenating the point position and feature of each point in the noisy point cloud and projecting them with a linear layer. Similarly, the $t$ token is obtained by projection with its own linear layer onto the same dimensionality. After encoding, all tokens are processed by transformer layers, including self-attention and MLPs. Finally, the resulting output tokens corresponding to the $M$ points are projected back to the dimensions of the input point positions and features and interpreted as noise predictions $\mathbf{\epsilon}^{\mathbf{P}}_{\theta}$ and $\mathbf{\epsilon}^{\mathbf{F}}_{\theta}$.
@@ 3.4 Disentangled generation
For appearance-only sampling from $p(\mathbf{F}|\mathbf{P})$ with given point positions $\mathbf{P}_{0}$, we obtain the initial noisy neural point cloud by sampling $\mathbf{F}_{T}$ from a unit Gaussian distribution and computing $\mathbf{P}_{T}$ from $\mathbf{P}_{0}$ via the forward process in Eq. 7.
Throughout the diffusion process, we update the point features $\mathbf{F}_{t-1}$ from the denoiser outputs according to the reverse diffusion process in Eq. 9, but update the point positions $\mathbf{P}_{t-1}$ from the given $\mathbf{P}_{0}$ via the forward diffusion process in Eq. 7 instead of the denoiser outputs. Sampling from $p(\mathbf{P}|\mathbf{F})$ is done analogously. Further details are provided in the supplementals.
Intuitively, this is comparable to masked image inpainting, using an approach similar to RePaint [26]. Instead of masking image parts, we mask one modality of our representation. Note that this is enabled by the neural point representation, which allows to disentangle the variables for coarse shape and local appearance.
@@ 4.1 Datasets and experimental setup
Additionally, we use the PhotoShape Chairs dataset [34]. The dataset contains $15{,}576$ objects and features more realistic textures on top of ShapeNet meshes. We use the same test split as DiffRF [30] with $1,552$ objects. From the remaining objects, we randomly select $2,480$ objects for training. We use the same renderings as DiffRF [30], which consist of 200 views per object on an Archimedean spiral. We use a resolution of 128x128 pixels and the same point clouds with 512 points as for SRN chairs.
@@
Figure 2: Qualitative examples of disentangled generation on SRN cars, SRN chairs, PhotoShape chairs. 2(a) Appearance-only generation: we show a generated object and objects with re-sampled appearance. 2(b) Shape-only generation: we show a generated object and objects with re-sampled coarse shape. We can get diverse samples of local appearance or coarse shape when the respective other is given.
@@
We construct neural point clouds by zero-initializing features for each point. In the Point-NeRF autodecoder, we optimize the reconstruction loss in Eq. 3 with the TV and KL regularizers from Eq. 4 and Eq. 5. Further details on network architectures and training parameters are provided in the supplementals. For the diffusion model training, we normalize the neural point clouds and use DDPM [16] diffusion model parameters. Further details on the denoiser architecture, diffusion model parameters, and training parameters are provided in the supplementals.
@@ 4.6 Analysis
As diffusion on hybrid point clouds and local radiance fields has not been done before, we conduct ablation studies and analyze various novel design choices. Here, we analyze the effects of different initialization strategies, feature dimensionality and regularization methods in the category-level Point-NeRF autodecoder and diffusion model. We provide a more detailed analysis in the supplementals.
@@
As we assume that the structure of the feature space strongly affects the diffusion model, we analyze the effects of applying KL and TV regularizations to the neural point features during the Point-NeRF autodecoder training. Tab. 4 shows that both regularizations further decrease the ambiguity of the latent space over the zero initialization. We can summarize that a combination of TV and KL regularization provides overall the best results (c.f. supplementals). Overall, we found an appropriate initialization and regularization to be key ingredients for successful neural point cloud diffusion.